OpenStudy (anonymous):

There are 100 gumdrops. 73 of them are large, and the rest are small. 39 of them are red, and the rest are other colors. 19 of the gumdrops are large red ones. You choose 1 gumdrop at random, and it is that yummy red color. What is the probability that the gumdrop is large, given that it is red? Write fractions using the slash ( / ) key. Reduce fractions to their lowest terms.

OpenStudy (anonymous):

let L be the event of the chosen gummy bear being Large and let R be the event of the chosen gummy bear being Red. We know P(L)=0.73 P(R)=0.39 and the tricky part, P(L intersection R)=0.19 we gotta find P(L|R)=P(L intersection R)/P(R)=0.19/0.39=19/39